If a constant torque of 500Nm turns a wheel of moment of inertia 100kgm² about an axis through its center find the gain in anguler velocity in 2 seconds
1
Expert's answer
2020-04-08T10:38:43-0400

J=500Nm - torque

I=100kgm² - moment of inertia

"\Delta"t=2s - time interval

"\alpha" - angular acceleration

"\Delta\omega" - ? - gain in anguler velocity

"J=I\alpha \\
\alpha=\frac{J}{I}\\
\alpha=\frac{500}{100}=5 \ rad/s^2\\
\Delta\omega=\alpha \Delta t\\
\Delta\omega=5*2=10\ rad/s"


Answer: "\Delta\omega=10\ rad/s"